Wavelength of the balmer Halpha line(first line) is 6565 angstrom.  Calculate the wavelength of Hbeta (second line).

Grade:11

1 Answers

Arun
5 years ago
Lyman series start from n1 = 1 and n2=2,3, 4 , 5,.....
so, second member of Lyman series 
n1 = 1 and n2=3 
now,
1/lemda =R.Z^2[1/n1^2 -1/n2^2]
=RZ^2[1/1^2 -1/3^2] ------(1)
now, 
given,
first member of balmer series 
1/6563 =R.Z^2[1/2^2 - 1/3^2] --------(2)
divide equation (1) and (2)
6563/lemda =(1-1/9)/(1/4-1/9)=(8/9)/(5/36)
=32/5
lemda =6563 × 5/32 =1025.48